OpenStudy (anonymous):

I don't know what to do with this problem. It's 3x-y^2=16 It wants to know if it's a function. I know what a function is, and how to figure out if an equation is a function, and why or why it may not be, but this one stumps me.

OpenStudy (anonymous):

of course it is not a function it is an equation

OpenStudy (alexandervonhumboldt2):

-y^2=16-3x y^2=3x-16 y=sqrt(3x-16)

OpenStudy (anonymous):

the real question is, if you solve for \(y\) what do you get lets do it carefully

OpenStudy (alexandervonhumboldt2):

this is a function @satellite73

OpenStudy (anonymous):

ur no

OpenStudy (alexandervonhumboldt2):

oh yeah

OpenStudy (anonymous):

\[ 3x-y^2=16 \]is an equation, it is not a function the person who wrote this question is confused, but what should really be asked is can \(y\) be function of \(x\) and the answer is NO

OpenStudy (anonymous):

\[3x-y^2=16\\ 3x-16=y^2\\ \pm\sqrt{3x-16}=y\] and the \(\pm\) tells you x is not a function of y

OpenStudy (anonymous):

or rather y is not a function of x

OpenStudy (anonymous):

Oh wait, I'm sorry, I forgot the 4. The equation is actually 3x-4y^2=16. @satellite73 Equations are functions. You plug in x values to get y values. An equation is a function if only one y value comes out per x value.

OpenStudy (anonymous):

right the 4 makes no difference the answer is still NO because of the square on the y

OpenStudy (anonymous):

The 4 makes a difference in the equation, though. So, what does the equation turn out to be when solving for y?

OpenStudy (anonymous):

ahhh thus begins the confusion between equations and functions functions CAN be given by equations, but they do not have to be and all equations are not functions

OpenStudy (anonymous):

\[ 3x-4y^2=16. \\ 3x-16=4y^2\\ \frac{3x-16}{4}=y^2\\ \pm\frac{\sqrt{3x-16}}{2}=y\]

OpenStudy (anonymous):

I know that, I didn't say all equations were, I just said that equations can be functions, too.

OpenStudy (anonymous):

in any case it is the \(\pm\) that throws it off be careful when solving with a square common mistake would be to write \[y=\frac{\sqrt{3x-16}}{2}\] and forget about the \(\pm\) part

OpenStudy (anonymous):

another way to see it is not a function is to know that this is a parabola that opens to the right, so does not pass the vertical line test http://www.wolframalpha.com/input/?i=3x-y^2%3D16+

OpenStudy (anonymous):

So, the question still remains, is this a function. Does every x input have only one y output?

OpenStudy (anonymous):

and the answer is NO we can do it by example if you like

OpenStudy (anonymous):

can we find two different y values that come from the same x value in other words, and the answer is yes we can pick an x and we can find the two y values

OpenStudy (anonymous):

I graphed it on my graphing calculator, and got half a sideways parabola, meaning it would pass the vertical line test.

OpenStudy (anonymous):

did you get something like this ?

OpenStudy (anonymous):

No, like I just said, I got half a sideways parabola.

OpenStudy (anonymous):

that is because your graphing calculator only graphs functions, so it probably graphed the top half only

OpenStudy (anonymous):

like i said, the square on the y is what makes it not a function just like \(x=y^2\) is not a function since if \(x=25\) \(y=5\) or \(y=-5\) both work